OpenStudy (anonymous):

find the funtion value f(x)=x3

OpenStudy (anonymous):

\[^{?}3\]

OpenStudy (anonymous):

x to the 3rd power

OpenStudy (zarkon):

x^3 \[x^3\]

OpenStudy (anonymous):

so what would be the answer 3?

OpenStudy (phi):

You can't do anything more unless you have a value for x otherwise it is just x^3

OpenStudy (anonymous):

8\[^{3}\]

OpenStudy (zarkon):

\[8^3=512\]

OpenStudy (phi):

If the question is for x=8: f(8)=8^3 = 8*8*8= 512

OpenStudy (phi):

try the equation editor, or just stick to uparrow to show exponentiation.